# Exploring Families’ Use and Mediation of Generative AI: A Multi-User Perspective

### v2

@@ 4.4.3. Mediation Strategies $\Leftrightarrow$ Perceived Opportunities & Risks of GenAI
5. Family Profiles: Instantiating the Family-GenAI Mediation Model
In this section, we elaborate on specific findings with family profiles that explain the Family-GenAI Mediation Model described in Section 4. These case studies provide practical examples of how the Family-GenAI Mediation Model is instantiated in practice. Table 3 represents a summary of notable characteristics of each type of family with respect to control, trust, and co-use patterns.
Family Profiles Description (Family ID)
A Skeptical Family Dimensions (F6): High control, low trust, medium co-use
Characteristics: Concerned toward the safety of GenAI; sees value in technology use; mediates concerns by setting protective measures for using ChatGPT
Representative Quote: “I’m putting some guardrails around which programs and apps my child uses.” (P6)
A Cautious Family Dimensions (F10): High control, high trust, low co-use
Characteristics: Concerned about exposure to developmentally inappropriate content; trusts the platform’s safety measures; mediates concerns by setting exploratory rules for monitoring ChatGPT use
Representative Quote: “I would like to keep an eye on them while my child is using ChatGPT” (P10)
A Hands-On Family Dimensions (F9): High control, medium trust, medium co-use
Characteristics: Concerned about ethical boundaries of cheating; trusts school policies and guardrails; mediates concerns by actively screening children’s ChatGPT use
Representative Quote: “My child comes to me. I use my personal judgment on if it’s appropriate or not to ask [ChatGPT something].” (P9)
A Use Together Family Dimensions (F5, 8, 11): High control, medium trust, high co-use
Characteristics: Concerned about ethical misuse and developmental risks for children; trusts the information quality or children’s self-regulation; mediates concerns by actively using ChatGPT together and requiring permission for use
Representative Quote: “[We] try to share accounts with [my child] so we can, you know, see what they use it for and so that we can talk about what they’ve used it for, things like that.” (P11)
A Trusting Family Dimensions (F2, 3, 7, 12): Medium control, high trust, medium co-use
Characteristics: Concerned about beyond-human capabilities of ChatGPT; trusts ChatGPT’s built-in protections; addresses concerns by co-using ChatGPT, but sometimes puts trust and responsibility in ChatGPT’s existing safeguards
Representative Quote: “I mean I don’t really have any [concerns]… There are limits built in[to ChatGPT]. So I think it’s one of the interesting tools that I actually feel [I can trust].” (P7)
An Independent Family Dimensions (F1, 4): Low control, medium trust, low co-use
Characteristics: Relaxed on rules
Representative Quote: “My child has to learn how to navigate and do this stuff himself. I can’t always be there for them.” (P1)
Table 3. Observed Family Profiles: We represent each category of case studies and describe how the Family-GenAI Mediation Model is instantiated for each observed family category, grouped by degree of control. We discuss the dimensions of each family with respect to the control, trust, and co-use patterns. We describe family characteristics and share a representative quote of the family categories.
5.1. A Skeptical Family: Sets protective measures for prompting ChatGPT
While the “Skeptical” family in this study (F6) acknowledged that using technology can sometimes be beneficial, they were wary about the safety of GenAI and utilized protective measures when using ChatGPT. Despite prior experience using GenAI-based products, P6 did not feel that they could trust ChatGPT: “ChatGPT is new, and I’m not sure how secure it is.” One of their main concerns with ChatGPT stemmed from their experiences with privacy breaches online; P6 stated that they had received many emails from companies about their data being leaked and decided to implement “some guardrails around which programs and apps [C6] uses to protect their child from having their personal information mishandled. This concern and practice transferred to other GenAI products. Although they mentioned trying GenAI-based online therapists, P6 said, “I do not give real data about myself”. Instead, they stated that they would “give it, like, generic scenarios, or I’m gonna fake a scenario. But I know what the emotional content is about in that scenario.” These concerns were then communicated to the child through conversations about ChatGPT and GenAI.
5.1.1. Suspicions for Location Tracking and GenAI
Both C6 and P6 distrusted what ChatGPT stated about its capabilities. During the interview, we encouraged the family to test whether ChatGPT actively tracks users’ location data. F6 asked ChatGPT to locate a restaurant near them, and ChatGPT consistently responded that it could not access real-time location data and recommended using online map applications instead. The response did not include any solicitation from the user to share their location, but both C6 and P6 felt that ChatGPT was lying. “It’s lying. And now I think it does [know our location],” C6 stated, with P6 adding, “I think it knows [our location]… but it’s not allowed to say because it would, hmm, violate [OpenAI’s] terms of service.” P6 also felt GenAI-based products were tailored to the user, so users had to be more careful about what they shared with GenAI. They explained that they would “lie to [Gen]AI”, but they would not “lie to Google”: “I use Google for health stuff. But I Google so many things that Google must be really confused about me … I might even be helping a friend because her child has a rash. So then, I say, ‘What does measles look like?’ But it’s not about me.”
5.1.2. Family-GenAI Co-use in Games
Despite their skepticism and distrust toward ChatGPT, the family also had positive co-use experiences on the platform. F6 shared that they often played games together but were limited to two-player games as they were the only people in their household. ChatGPT’s interactive, human-like features enabled new gameplay structures for F6. P6 said, “You know, it’s hard to play a lot of games with just two people. So ChatGPT plays that [role]. They’re holding the cards… So then that helps us play a good game.” They also encouraged C6 to interact with ChatGPT: “I really believe that technology is so useful that it would be harmful for me to keep [C6] off of his screens, because he’s learning and leaping forward massively.” Acknowledging the potential risks and benefits of ChatGPT, P6 decided to encourage their child’s use while still setting protective restrictions. During the interview, C6 often validated P6’s feelings and reasoning for setting the rules that they did. The “Skeptical” family frequently discussed risks and opportunities of GenAI: “We’ve talked at dinner and at lunch… stuff about, ‘do we have any concerns about the usage of [Gen]AI?’, and we both do have some concerns about it.”
5.2. A Cautious Family: Sets exploratory rules for monitoring ChatGPT use
The “Cautious” family in this study (F10) trusted ChatGPT’s in-platform safeguards and safety features, but P10 still monitored C10’s use because they were worried that ChatGPT might generate inappropriate output. Having self-disclosed as being proficient in programming and technology, P10 explained that they knew ChatGPT had safety mechanisms. However, they also mentioned that they knew these mechanisms could be bypassed: “ Earlier this year, [there were] some kind of trick prompts [people used to dupe ChatGPT successfully].” Thus, they still felt it was necessary to monitor their child’s use of ChatGPT closely: “I know technology always has some [issues]… I would like to keep an eye on [my children] while [they are] using ChatGPT.” Relatedly, P10 felt that it was important for them to understand any technology they allowed their child to use.
5.2.1. Distinct Mediation of ChatGPT and Social Media
While P10 actively monitored C10’s ChatGPT interactions after they occurred (and were thus labeled as having high control over their usage), they did not transfer some of their rules for other technologies to GenAI. For example, C10 was allowed to use age-restricted Google and YouTube accounts and a personal device with a child lock on it, but P10 was not concerned about having this type of mediation for ChatGPT. Although this may be due to the lack of age-restricted accounts on ChatGPT, P10 felt that ChatGPT was not “addictive”, but social media could be (e.g.,  Al-Samarraie et al. (2021)). When discussing why they disliked social media, they stated, “I think there’s a difference… ChatGPT [gives you] the information. It’s not necessarily [playing] those sorts of [addictive] video… [or] a quick song like YouTube.” While discussing social media, P10 said, “You get nothing from it, just feel lost. And there’s lots of studies showing those short clip videos has [sic] a very negative impact on your- on your brain, on your mental health, and addiction… Although social media is just to have a very tiny bit of information… It’s- it’s not an article. An article gives you a way to really think those things through. If you want to explain something, want to present something, you-you’d… better read an article instead of 140 characters on Twitter, or [watching] those 20-second videos.”
5.2.2. Contextual Factors: Family-GenAI Communications around GenAI and Rule Exploring.
Besides monitoring C10’s interactions with GenAI, P10 had conversations about the technology with them to build their knowledge about it. F10 did not co-use ChatGPT, as P10 felt it was important for children to explore the technology on their own (with monitoring) to understand how it worked: “If you make the rule, you need to make sure [the child] understands the rule. If you understand something, it becomes your knowledge.” As far as whether they had plans to implement any usage rules for ChatGPT, P10 said, “ I don’t have any- any kind of a [specific ChatGPT] rule… I trust it. For now.”
5.3. A Hands-On Family: Actively screens child’s ChatGPT use
The “Hands-on” family (F9) actively screened their children’s ChatGPT questions or prompts either before or after they were submitted to the platform. Pa9 and Pb9 were mainly concerned about whether their children could clearly distinguish between using ChatGPT to cheat versus to support learning. Pb9 aimed to make sure that their children “don’t just copy and paste” or plagiarize ChatGPT outputs for their own use. As a result, they required their children to ask them for permission to use the platform and reviewed the types of prompts their children wanted to input before doing so: “They come to me. I use my personal judgment on if it’s appropriate or not to ask.”
5.3.1. Parents’ Prioritize Traditional Resources over ChatGPT
Both parents believed that although learning how to use ChatGPT might benefit their children’s futures (e.g., if they need to use it at a job), it was not a necessary tool in schools, with Pb9 stating, “except for research[ing information], there’s really no need for ChatGPT.” Both parents preferred that their children use other resources, such as Google, and encouraged their children to approach them for help. In characterizing themselves as trustworthy sources of information, Pb9 dubbed themselves and Pa9 as the “older ChatGPT”, or a more rudimentary form of ChatGPT.
5.3.2. Family-GenAI Creative Co-use Patterns
Despite their concerns about ChatGPT, both parents actively encouraged their children to interact with the GenAI to learn about what it could do. Having learned about ChatGPT through their technology-based job, Pb9 decided to create a shared account “just for educational purposes only … [to] showcase what’s possible for the kids.” Pb9 would occasionally co-use ChatGPT with C9 and S9 as a writing tool (e.g., to edit an essay or explore a story idea). Pb9 mentioned that although they wanted to be “mindful” about risks, they also did not want to “restrict [C9 and S9] from access [to] being creative” online.
5.3.3. Contextual Factors: Challenges in Managing ChatGPT Use at School
F9 had general parental controls on the children’s devices and router-level settings configured that filtered through content on all platforms, including ChatGPT. However, they highlighted that these mediation strategies were not foolproof. For instance, Pb9 hoped that the school district would set up some guardrails for ChatGPT use, but acknowledged that, “It’s a- a challenge in general… Unless you build a special custom portal that has all these guardrails around what you can or cannot ask.”
5.4. A Use Together Family: Actively co-using ChatGPT together
“Use Together” families (F5, F8, F11) frequently co-used ChatGPT with one another. All three families shared accounts, and the children in every family reported minimal instances of using ChatGPT alone (sometimes due to a lack of desire to do so) and mentioned that they had to receive permission from their parents to use it. None of the children were interested in opening a personal ChatGPT account. Additionally, none of the participating children in these families had personal devices
5.4.1. Contextual Factors: Occupations, Concerns, and Family Conversations
A distinguishing factor of “Use Together” families was their perceptions of ChatGPT. For these three families, parents’ concerns were driven by what they knew about ChatGPT from their jobs. For instance, both P8 and Pb11 were teachers, and both mentioned having students who used ChatGPT to cheat. Furthermore, most of the parents (besides Pb11) self-described themselves as not familiar with how ChatGPT functioned. P5 and P8 both explicitly expressed confusion about how ChatGPT collects user data, and Pa11 mentioned that as someone who was not “tech-savvy”, they were nervous about what ChatGPT could look like in the future. P5, who was also concerned about data safety, said, “If [OpenAI or ChatGPT] are ever hacked… [Hackers would] know that [a] person at this address has three daughters.” P5 was unaware of some of the platform’s safety measures; when they asked ChatGPT to identify a good bakery, they found it odd that it recommended asking a local friend for an answer instead. P8 was confused about ChatGPT’s data collection practices and transparency: “In terms of thinking about the data that it collects… I don’t understand fully what ChatGPT is taking from me.” Despite these confusions and concerns, the “Use Together” families expressed that they feel safe using ChatGPT due to their harmless use cases. P5 and P11 both trusted their children to use ChatGPT safely. P8 considered ChatGPT an entertaining tool for the family: “So it’s still very much like more creative fun. Silly at home [stuff].” P11 also explained that discussing professional concerns about students cheating and online safety with their children helped raise awareness of the risks and opportunities of using ChatGPT: “I think she’s heard me complain enough about the threat of students cheating and how- how obvious it is when students plagiarize, especially from things like ChatGPT. So I think she kinda knows not to do that stuff. She’s kind of in and around these conversations about internet safety.”
5.4.2. Comparisons Between ChatGPT and Other Media
“Use Together” families perceived ChatGPT as safer than social media platforms, such as YouTube. Thus, parents had more relaxed rules for ChatGPT compared to other platforms. P8 mentioned that they knew “a lot of researchers who have talked about the dangers [of social media], like the psychological issues behind young people with cell phones and having access to social media.” They also pointed out that C8 did not have access to any social media accounts and was not allowed to post public comments on YouTube. Pb11 shared similar concerns about social media, but focused more on the content presented on the platforms. They considered text-based LLMs like ChatGPT to be safer than social media because no visual content was being presented to users: “I think because ChatGPT has certain features that are just very different from those other kinds of social media platforms, then it seems a little safer in that way. I mean, there’s obviously still risks and things that she could get into that might be nefarious, too, but at least [ChatGPT] doesn’t have some of those qualities that these other [social media] platforms have.”
5.4.3. Family-GenAI Co-Use Patterns
Compared to the other families in this study, “Use Together” families reported frequent co-use of devices and media, including ChatGPT, and minimal rules specific to ChatGPT. Parents in this group felt that co-use could enhance bonds between family members. For example, F5 regularly worked on math problems from school together and had begun using ChatGPT to review and discuss their answers. F5 also mentioned that they sometimes used ChatGPT to help resolve family conflicts: P5 brought up a time they asked ChatGPT to explain the steps in solving a math problem, as P5 and C5 disagreed on the ‘proper’ way of doing so. F8 and F11 also mentioned co-use cases where the family used ChatGPT to generate stories (e.g., short stories, fanfiction) or written work together.
5.5. A Trusting Family: Transfers responsibility to ChatGPT to support mediation
“Trusting” families (F2, F3, F7, F12) had few concerns specific to GenAI and were trusting of GenAI, but still employed several online safety rules and sometimes applied mediation strategies together as a family. P7 and P12 had minor concerns, while P2 and P3 had no concerns about their children’s ChatGPT use. Both P7 and P12 were educators, which contributed to their worries about their children cheating and over-relying on ChatGPT. In describing their concerns, P12 said, “I think there’s the academic integrity impacts. And then, just baseline writing ability and creativity stunting that might happen if you were to rely entirely on these models than to kind of create any sort of academic outputs that you’re being asked to do.” P7 was also nervous that GenAI use might discourage creativity: “…And then [Gen]AI generated those four things, and then [my child] put those four things on his poster… What could happen [in the future]? It’s all fancy, and it took me, you know, five minutes to [generate] four things [on ChatGPT]. And this is why [GenAI] is problematic- because it takes away human creativity.”
5.5.1. Reasons for Trust in GenAI
Two families (F2, F7) discussed why they had high trust in ChatGPT and minimal safety concerns about the platform. F2 described that they trusted ChatGPT over Google because it had protective measures. To demonstrate this during the interview, C2 tested both ChatGPT and Google by inputting P2’s phone number into the prompt box. ChatGPT did not return anything, while Google provided detailed personal information based on the number. P2 was impressed by ChatGPT’s response but was worried by Google’s: “When he Googles, it gives my age, my address… So yeah, [my child’s other parent] comes up first, which is weird.” Notably, P2 had limited knowledge about how ChatGPT is trained from past user data. In terms of the risks of using ChatGPT, P7 only cited minor concerns, and said, I don’t really have [concerns]… to me, it seems like, well, there are limits built in[to ChatGPT]. So I think it’s one of the interesting tools that I actually feel [I can trust].”
P2, P3, and P12 trusted their children to use ChatGPT safely and appropriately. Although P12 did not have particularly high trust in ChatGPT, they trusted C12 to make smart decisions online, stating, Yeah, we are operating with a lot of trust [in our child]. [She] is a good kid who’s never given us any reason to doubt.” C12 then mentioned that they only used ChatGPT to generate images when using the platform alone; F12 did not mention any concerns about ChatGPT outputting inappropriate imagery. Similarly, P2 trusted C2 and ChatGPT’s safeguards, stating, “He would have to type in something very specific to get to something inappropriate on ChatGPT… I trust him that he’s not looking up things he shouldn’t be.” P3 did not feel that C3 had any nefarious reasons to use ChatGPT, stating, “I mean, that might come a point where I, you know, I don’t like what he’s doing with [ChatGPT]. But right now, I think it helps his language and vocabulary, and he’s doing something creative with it.”
5.5.2. Family-GenAI Co-Use and Communications around GenAI
“Trusting” families often discussed GenAI and online safety with one another and sometimes co-used ChatGPT, including exploring the mechanics of the technology together (F7) and generating content together (F7, F12). When explaining why they proactively learned about ChatGPT together, P7 said, “…With the Internet, we’re teaching them how to do those [appropriate] things and how not to do those [inappropriate] things.” C7 and S7 would also generate stories based on their parents’ occupations and share parts of the output with their parents. P12 highlighted a positive experience co-using ChatGPT with C12 to generate a story: “We were like, ‘write a story about an obese horse who is working to lose some weight.’ And it came up with this amazing story. I think we even made it so that the horse’s name was going to be Charger, just like our obese horse, and it was a great story about this horse who, you know, was gonna stay away from the sugar cubes and focus more on eating vegetables and grass and exercising. And… what a great story!”
5.6. An Independent Family: Relaxed on rules
@@
In “Independent” families (F1, F4), each family member used ChatGPT individually. Parents did not enforce any controls, and both families were open with each other about their technology use. For example, C1’s technology use (including ChatGPT) was rarely controlled at home, and the only mediation strategy that P1 employed was checking C1’s device for any exposure to inappropriate content from other people: “It’s more the people he’s talking to. So I just wanna make sure that they’re being appropriate… I trust him. I don’t trust other people.” P4 did not have family rules: “We’re quite lax that way.” P1 exhibited minimal concerns about GenAI and expressed trust in C1’s usage: “Oh, no, I mean, he’s a good kid, and we taught him to tell the truth. So I’m not really concerned about it. He knows right from wrong.” P1 also felt that children should learn about technologies independently because they would eventually have to navigate them on their own: “He has to learn how to navigate and do this stuff himself. I can’t always be there for him.” P4 did not feel concerned about their children asking ChatGPT inappropriate questions. However, they paused the study at one point to teach their children about “cross-validating information” and “researching different information sources” by comparing the strengths and limitations of different technologies (e.g., Google, Alexa, smartwatches). P4 also indicated that children nowadays consider GenAI their “search engine” by comparing prompt engineering and searching keywords: “Whereas I am still old school. I was there when the first search engines were made, and we were taught how to phrase the terms in order to get the best results.” Both F1 and F4 discussed technology use with their children often, which contributed to their relaxed attitude toward rules. P1 stated, “We don’t really have any [secrets]… Everybody knows the password of everything.”

### v3

@@ 4.2.4. Connections between Factors
4.3. Revisiting Families Through the Lens of the Family-GenAI Mediation Model
In what follows, we describe each family’s interactions around GenAI through the lens of our proposed Family-GenAI Mediation Model. We also present a detailed summary of how components of the Family-GenAI Mediation Model are distributed in families in Table 2.
4.3.1. Family 1: No Mediation Strategy
@@
F1 did not have any rules around children’s GenAI usage. Even for children’s general online safety, the only general online safety mediation strategy that P1 implemented was periodically checking C1’s device for any exposure to inappropriate content from other people. As P1 explained, “It’s more the people he’s talking to. So I just wanna make sure that they’re being appropriate… I trust him. I don’t trust other people.” P1 did not hold this concern toward ChatGPT, and they did not consider ChatGPT to be directly beneficial to their children. “I just normally use it just to get information.”, P1 concluded. Regarding their children’s ChatGPT usage, P1 exhibited minimal concern over GenAI and expressed trust in C1: “I mean, he’s a good kid, and we taught him to tell the truth. So I’m not really concerned about it. He knows right from wrong.” P1 also felt that children should learn about technologies independently because they would eventually have to navigate those technologies on their own: “He has to learn how to navigate and do this stuff himself. I can’t always be there for him.” F1 often has open communication and in-family discussions about technology use, which contributed to their relaxed attitude toward rules. P1 stated, “We don’t really have any [secrets]… Everybody knows the password of everything.” This sentiment illustrates P1’s perception that transparency in technology use can be used in place of strict controls or interventions.
4.3.2. Families 2, 3, & 7: Co-use & Supervision
F2, F3, and F7 all employed both co-use and supervision mediation over children’s GenAI usage. Their children needed to access ChatGPT over a shared device, and they sometimes shared the GenAI-generated content with their parents. While C7 and S7 have individual devices, they normally access ChatGPT on the shared family computer, as their school has restrictions on GenAI usage. In F2 and F3, parents’ mediation strategies mirror their preexisting rules and expectations for general online safety, namely monitoring children’s device usage.
These parents also held a positive attitude toward their children’s use of ChatGPT and did not have any concerns about GenAI. P7, an educator, however, is worried about their children cheating or over-relying on ChatGPT. P7 was particularly nervous that GenAI use might discourage creativity. “…And then [Gen]AI generated those four things, and then [my child] put those four things on his poster… What could happen [in the future]? It’s all fancy, and it took me, you know, five minutes to [generate] four things [on ChatGPT]. And this is why [GenAI] is problematic- because it takes away human creativity,” P7 stated.
In terms of the risks of using ChatGPT, P7 only cited minor concerns and said, “I don’t really have [concerns]… to me, it seems like, well, there are limits built in[to ChatGPT]. So I think it’s one of the interesting tools that I actually feel [I can trust].” P3 also considered ChatGPT to be safe because it is an AI product: “It’s just him and the AI interacting, so that I don’t really have privacy concerns so much cause he’s not actually talking to a person.” F2 similarly trusted ChatGPT more than Google due to its protective measures. This rationale was demonstrated during the interview when C2 input P2’s phone number into both platforms; while ChatGPT returned nothing, Google revealed personal details. This example shows that co-use of GenAI can shape families’ perceptions and, in turn, influence the manner in which they mediate children’s use of these platforms.
Despite minor concerns, P2 and P3 trusted their children to use ChatGPT safely and appropriately. As P2 stated, “He would have to type in something very specific to get to something inappropriate on ChatGPT… I trust him that he’s not looking up things he shouldn’t be.” P3 did not feel that C3 had any nefarious reasons to use ChatGPT, and considered ChatGPT to be beneficial as an educational and creative tool. This sentiment was illustrated by P3 who stated, “I mean, there might come a point where I, you know, I don’t like what he’s doing with [ChatGPT]. But right now, I think it helps his language and vocabulary, and he’s doing something creative with it.” We also observed that P2 and C2 had open communication regarding GenAI use. For instance, C2 admitted honestly that they created an account in school to look up math problems without informing P2 because they were not allowed to do so. Then P2 praised C2’s honesty, stating, “Oh, good! (touching C2’s face) Cheating! That’s good. (laugh)”. Additionally, P7 specifically considered that they should teach children to use new technologies instead of restricting them. P7 said, “…With the Internet, we’re teaching them how to do those [appropriate] things and how not to do those [inappropriate] things.” Furthermore, these families considered co-use an interesting way to support family interaction. C3 reported normally using ChatGPT to create scripts for YouTube videos and to share those videos with family members. C7 and S7 would also generate stories based on their parents’ occupations and share parts of the output with their parents. P2 mentioned an interesting case where they solved a family debate by asking ChatGPT: “My husband and I are in agreement that it’s called the Spikes Shell. But yeah, he and his brothers call it a blue shell… [we] just asked it if that’s a blue shell.”
4.3.3. Family 4: Instructive Mediation
The only mediation F4 utilized was instructively communicating with children about GenAI. P4 exhibited high familiarity towards GenAI: “I was involved in a project that helped train Bard.” While P4 expressed many concerns (e.g., concerns regarding misinformation and user privacy), they reported a nonchalant attitude toward technology use in their home: “No [rules]. We’re quite lax that way.” Instead, P4 stated they encourage their children to access and use new technologies and teach them the appropriate way to navigate those technologies. During the interview, P4 at one point paused the study to teach their children about “cross-validating information” and “researching different information sources” by comparing the strengths and limitations of different technologies (e.g., Google, Alexa, or smartwatches). P4 also noted that children of the present generation consider GenAI to be their “search engine” by contrast to P4’s prior experience with actual search engines. P4 reflected, “… I am still old school. I was there when the first search engines were made, and we were taught how to phrase the terms in order to get the best results.” While the experience of P4 drastically differs from that of their children, P4’s knowledge of past and present technologies provides them with insights that they may use to instruct their children on GenAI use.
4.3.4. Families 5, 8 & 11: Restrictive Mediation & Co-use
F5, F8, and F11 utilized both co-use and restrictive mediation. One commonality all of these families share is using ChatGPT together as a family. Additionally, these families have established rules for online safety that include the children not having their own individual devices. As a result, the children in these families must first ask their parents before they can use ChatGPT. In terms of risks and benefits, parents usually felt there was a trade-off in using GenAI. For example, P5 reported a lack of familiarity with GenAI, and relatedly expressed concerns, such as about ChatGPT data collection and data safety policies. As P5 stated, “If [OpenAI or ChatGPT] are ever hacked… [hackers would] know that [a] person at this address has three daughters.” P5 was unaware of some of the platform’s safety measures; when they asked ChatGPT to identify a good bakery, they found it odd that it recommended asking a local friend for an answer instead. F8 and F11, however, held a different set of concerns towards GenAI. Due to P8 and P11b’s academic occupations, they were more concerned about the ethical use than the data safety. Despite that fact, both parents reported feeling comfortable with their children’s use of ChatGPT and did not specifically set any rules around these concerns. F8 reported using ChatGPT for creative use cases: “So it’s still very much like more creative fun. Silly at home [stuff].” P11b expressed trust in their child to use GenAI appropriately: “I think she’s heard me complain enough about the threat of students cheating and how- how obvious it is when students plagiarize, especially from things like ChatGPT. So I think she kinda knows not to do that stuff. She’s kind of in and around these conversations about internet safety.” P8 and P11b also considered ChatGPT to be safe in terms of content, because only minor visual content (e.g., images and videos) and text are presented to users. P11b shared their opinion by comparing social media and GenAI: “I think because ChatGPT has certain features that are just very different from those other kinds of social media platforms, then it seems a little safer in that way. I mean, there’s obviously still risks and things that she could get into that might be nefarious, too, but at least [ChatGPT] doesn’t have some of those qualities that these other [social media] platforms have.” P8 also expressed similar concerns: “a lot of researchers … have talked about the dangers [of social media], like the psychological issues behind young people with cell phones and having access to social media.” P8 explained that C8 did not have access to any social media accounts and was not allowed to post public comments on YouTube. While these sentiments generally illustrate a greater concern with social media than GenAI, they ultimately showcase these families restrictive mindsets when mediating their children’s technological use.
Furthermore, the three families considered co-use of ChatGPT to be an opportunity to enhance the bonds between family members. For example, F5 regularly worked on school math problems together and had begun using ChatGPT to review and discuss their answers. F5 described an instance in which they asked ChatGPT to explain the steps involved in solving a math problem, as P5 and C5 disagreed on the ‘proper’ way of doing so. P8 recalled that they have used ChatGPT to generate fanfiction together: “Oh, we had ChatGPT write a poem about Jenny Weasley and Harry Potter’s love story.” F11 also mentioned co-use cases where the family used ChatGPT to generate stories (e.g., short stories, fanfiction) or other written work together.
4.3.5. Family 6: Instructive & Co-use & Restrictive
F6 employed various mediation strategies, including instructive mediation (e.g., having rules and conversations for using ChatGPT), restrictive (e.g., the child needs their parents’ permission to access ChatGPT), and co-use (e.g., the child only uses ChatGPT with the parent). Despite prior experience using GenAI-based products, P6 did not feel that they could trust ChatGPT: “ChatGPT is new, and I’m not sure how secure it is.” One of their main concerns with ChatGPT stemmed from their experiences with privacy breaches online; P6 stated that they had received many emails from companies about their data being leaked and decided to implement “some guardrails around which programs and apps [C6] uses” to protect their child from having their personal information mishandled. This concern and practice transferred to other GenAI products. Although they mentioned trying GenAI-based online therapists, P6 said, “I do not give real data about myself”. Instead, they stated that they would “give it, like, generic scenarios, or I’m gonna fake a scenario. But I know what the emotional content is about in that scenario.” These concerns were then communicated to their child through conversations about ChatGPT and GenAI.
Despite this distrust toward ChatGPT, the family also had positive co-use experiences. F6 shared that they often played games together, but were limited to two-player games as they were the only people in their household. ChatGPT’s interactive, human-like features enabled new gameplay structures for F6. P6 said, “You know, it’s hard to play a lot of games with just two people. So ChatGPT plays that [role]. They’re holding the cards… So then that helps us play a good game.” They also encouraged C6 to interact with ChatGPT: “I really believe that technology is so useful that it would be harmful for me to keep [C6] off of his screens, because he’s learning and leaping forward massively.” While F6 considered technology use to be beneficial at times, they were wary about the safety of GenAI and thus utilized protective measures when using ChatGPT.
4.3.6. Family 9: Co-use & Supervision & Restrictive
F9 exhibited multiple types of mediation: C9 needed to ask their parents’ permission to use ChatGPT (restrictive), share ChatGPT-generated content with parents (co-use), and have questions or prompts actively screened by their parents (supervision). P9a and P9b were mainly concerned about whether their children could clearly distinguish between using ChatGPT to cheat versus support learning. P9b aimed to make sure that their children “don’t just copy and paste” or plagiarize ChatGPT outputs for their own use. As a result, they required their children to ask them for permission to use the platform and reviewed the types of prompts their children wanted to input before doing so: “They come to me. I use my personal judgment on if it’s appropriate or not to ask.” The parents also encouraged their children to approach them for help. In characterizing themselves as trustworthy sources of information, P9b dubbed themselves and P9a as the “older ChatGPT”, or a more rudimentary form of ChatGPT. Despite the concerns, both parents actively encouraged their children to interact with GenAI to learn about what it could do. P9b decided to create a shared account “just for educational purposes only … [to] showcase what’s possible for the kids.” P9b would occasionally co-use ChatGPT with C9 and S9 as a writing tool (e.g., to edit an essay or explore a story idea). P9b mentioned that although they wanted to be “mindful” about risks, they also did not want to “restrict [C9 and S9] from access [to] being creative” online.
4.3.7. Family 10: Instructive & Supervision & Restrictive
F10 treated GenAI usage cautiously. P10 monitored C10’s use of ChatGPT (supervision), and C10 needed to ask for permission to use ChatGPT (restrictive). Additionally, F10 had family conversations about GenAI (instructive). The main concern highlighted by P10 is that ChatGPT might generate inappropriate outputs. Having self-disclosed as being proficient in programming and technology, P10 explained that they knew ChatGPT had safety mechanisms. They also mentioned, however, that they knew these mechanisms could be bypassed: “ Earlier this year, [there were] some kind of trick prompts.” Thus, they still felt it was necessary to monitor their child’s use of ChatGPT closely: “I know technology always has some [issues]… I would like to keep an eye on [my children] while [they are] using ChatGPT.” While P10 actively monitored C10’s ChatGPT interactions after they occurred, and were thus labeled as having high control over their usage, they did not transfer some of their rules for other technologies to GenAI. For example, C10 was allowed to use age-restricted Google and YouTube accounts and a personal device with a child lock on it, but P10 was not concerned about having this type of mediation for ChatGPT. Although this may be due to the lack of age-restricted accounts on ChatGPT, P10 felt that ChatGPT was not “addictive” like social media could be (e.g.,  al2021young). When discussing why they disliked social media, they stated, “I think there’s a difference… ChatGPT [gives you] the information. It’s not necessarily [playing] those sorts of [addictive] video… [or] a quick song like YouTube.” Besides monitoring C10’s interactions with GenAI, P10 also had conversations with them about technology to build their knowledge. F10 did not co-use ChatGPT, as P10 felt it was important for children to explore the technology on their own (with monitoring) to understand how it worked: “If you make the rule, you need to make sure [the child] understands the rule. If you understand something, it becomes your knowledge.” Regarding plans to implement usage rules for ChatGPT, P10 said, “ I don’t have any- any kind of a [specific ChatGPT] rule… I trust it. For now.”
4.3.8. Family 12: Co-use Only
The only mediation strategy F12 utilized for their child’s GenAI usage was co-use. Unlike other participating families, F12 shared a premium ChatGPT account for better utility and image generation. They also shared a social media account, while most families did not allow their children to have one. P12 felt it was safer for C12 to use ChatGPT than other online platforms: “The ability to connect in certain ways with people who may have… malicious intent. I guess those would be the biggest concerns.” The main concern of P12, like other parents with academic occupations, was children’s ethical use and over-reliance on GenAI. However, P12 did describe their trust and open communication with the family: “Yeah, we are operating with a lot of trust [in our child]. [She] is a good kid who’s never given us any reason to doubt.” Both C12 and P12 indicated that they would check on each other’s chat history for fun, and they would generate stories together as a family activity. C12 also brought up a scenario using ChatGPT with siblings: “Well, I hear some, like… my little sister could say something and then I’m like, ”Oh, I’ll put that into ChatGPT,” and then, yeah, just show it to them.”